BUSI 320 Comprehensive Problem 2 Spring 2019 You have been asked to assess the expected financial impact of each of the following proposals to improve the profitability of credit sales made by your company. Each proposal is independent of the other. Answer all questions. Showing your work may earn you partial credit. Proposal #1 would extend trade credit to some customers that previously have been denied credit because they were considered poor risks. Sales are projected to increase by $120,000 per year if credit is extended to these new customers. Of the new accounts receivable generated, 6% are projected to be uncollectible. Additional collection costs are projected to be 5% of incremental sales, and production and selling costs are projected to be 80% of sales. Your firm expects to pay a total of 30% of its income after expenses in taxes. 1) Compute the incremental income after taxes that would result from these projections: 2) Compute the incremental Return on Sales if these new credit customers are accepted: If the receivable turnover ratio is expected to be 4 to 1 and no other asset buildup is needed to serve the new customers… 3) Compute the additional investment in Accounts Receivable 4) Compute the incremental Return on New Investment 5) If your company requires a 20% Rate of Return on Investment for all proposals, do the numbers suggest that trade credit should be extended to these new customers? Explain. Proposal #2 would establish local collection centers throughout the region to decrease the time it takes to convert credit payments that are mailed in by check to cash. It is estimated that establishing these collection centers would reduce the average collection time by 2 days. 1) If the company currently averages $25,000 in collections per day, how many dollars will this suggested cash management system frees up? 2) If all freed up dollars would be used to pay down debt that has an interest rate of 8%, how much money could be saved each year in interest expense? 3) Do the numbers suggest that this new system should be implemented if its total annual cost is $7200? Explain.
Answer #1
Proposal # 1
Dollars Part 2 Working:
Incremental Sales 120000 Incremental Return on Sales = Incremental Income After Taxes / Net Sales
Collection Cost (5% of Incremental Sales) -6000 Receivable turnover ratio is expected to be 4 to 1 i.e Net Credit Sales / Account Receivables
Bad Debts (6% of Account Receivables (We assume Account Receivables as Incremental Sales i.e. all Credit Sales) -7200
Production & Selling Cost (80% of Incremental Sales) -96000 We are assuming Account Receivable to be 120000 then Net Credit Sales should be 120000*4=480000
Incremental Income Before Taxes 10800
Taxes @ 30% of Incremental Income -3240 Thus Incremental Return to Sales should be = 7560/480000 in %
Incremental Income after Taxes 7560 Part 1 Answer Part 2 Answer 1.58%
Part 3 Working:
To find the additional investment in Account Receivables
We know the Receivable Turnover Ratio is 4 to 1 and incremental sales is 120000
So Receivable Turnover Ratio in Days would be = 360/4 = 90 days it means that 90 days incremental sales are on average held up as receivables.
Part 3 Answer: = Additional Investment in Account Receivables = Per day incremental sales i.e. 120000/360 = 333.33 multiply by 90 days will give 30000
30000
Part 4 Working:
To find incremental return on New Investment we would use this formula = Incremental Income after Taxes / Incremental Investment Made in Account Receivable
Incremental Income after Taxes 7560
Incremental Investment 30000
Return on Incremental Investment 25.20% Answer Part 4
Part 5 Answer: Trade Credit can be extended to these customers as Return on Investment is 25.20% which is greater than 20.00%